# Efficient detection of chaos through the computation of the Generalized Alignment Index (GALI) by the multi-particle method

### v1

@@ 3.0.1 The variational method
In this section, we focus on evaluating the numerical uncertainty of the computation of GALI using the variational approach. It is worth noting that the VM constitutes the sole approach considered in previous studies (see e.g. [14, 15]) for the computation of the GALI. Thus later, we will prove that the definition of the GALI in the variational framework is equivalent to the definition using the MPM. The variational equations are obtained by linearizing the equations of motion [Eq. (2)] around a reference orbit $\bm{x}(t)$ in the phase space. Considering a set of $k$ orthogonal deviation (tangent) vectors
@@
where $\bm{D}^{2}\mathcal{H}\left(\bm{x}(t)\right)$ is the Hessian matrix of the Hamiltonian evaluated at the reference orbit $\bm{x}(t)$, and the subscript $i$ indexes the tangent vector. As initial conditions for the variational equations, the common practice is to choose random coordinates for the deviation vectors, which are then rescaled in such as way that they all share the same norm, i.e. $\lVert\bm{w}_{i}(0)\rVert=\lVert\bm{w}_{i,0}\rVert=d_{0}$ ($d_{0}>0$).
@@
In what follows, we first assume that the trajectory of the system $\bm{x}(t)$ and its tangent dynamics $\bm{w}_{i}(t)$ are known in a closed form. Consequently, we define the GALI of order $k$ (denoted as $G_{k}$ for the VM) of the orbit $\bm{x}$ after a renormalization time $\tau$ [14, 15]
@@
is the unit vector of the $i^{\textit{th}}$ deviation vector $\bm{w}_{i}(t)$.
We first investigate the error due to the machine precision (referred to as $\varepsilon$) during the computation of GALI after one renormalization time $\tau$ denoted as $G_{k}(\tau,\varepsilon)$. Starting from unit deviation vectors, i.e. $d_{0}=1$ and assuming that the renormalization time $\tau$ is small, we have $\lVert\bm{w}_{i}(\tau)\rVert\approx 1$ such that
@@
with $\widetilde{u}_{i,m}$ denoting the machine representation of $u_{i,m}$. In general, we will denote by $\widetilde{Q}$ any approximation of the variable $Q$. Thus the computer representation of the GALI of order $k$ writes
@@
in such as way that the error due to the machine precision writes
@@
In addition to effect of the machine precision on the computation of GALI, the magnitude of the initial deviation vectors $d_{0}=\lVert\bm{w}_{i}(t_{0})\rVert$ also impacts the computed GALI values at each renormalization time. By denoting $\widetilde{G}_{k}\left(\tau,\rho\right)$ the GALI obtained by replacing $\bm{w}_{i}(t_{0})$ by $\bm{w}_{i}^{\rho}(t_{0})=\rho\cdot\lVert\bm{w}_{i}(t_{0})\rVert$ where $\rho=d_{0}/\lVert\bm{w}_{i}(t_{0})\rVert$ is the rescaling factor, we see that the time evolution of the rescaled deviation vectors $\bm{w}_{i}^{\rho}(t)$ is given by
@@
and the magnitude of the $k$ initial deviation vectors do not affect the calculations of GALI in case of the VM. In this context, we only have to take into account round off errors for the computer representation of $\bm{u}_{i}^{\rho}$, similar to what was done in Eq. (13). Thus the numerical inaccuracy of the GALI at the renormalization time due to the magnitude of the deviation vectors leads to
@@
3.1 Multi-particle method
@@
the unit vector of the $i^{\textit{th}}$ separation vector in the phase space. Our first task here is to show that Eq. (20) is equivalent to Eq. (11) i.e. $J_{k}(\tau)=G_{k}(\tau)$, so that the GALI defined using the VM and the MPM are equal.
@@
to approximate the time evolution of $\bm{d}_{i}$ with $\mathcal{O}\left(\lVert\bm{d}_{i}\rVert^{3}\right)\rightarrow 0$, because the $\lVert\bm{d}_{i}\rVert\ll 1$. At this point, it is convenient for us to approximate the evolution of each element $d_{i,m}$ of the separation vectors with index $i$ by the dynamics of its fastest changing coordinate corresponding to the direction of maximal stretching of the tangent vectors with rate of growth being the MLE [7]. Thus Eq. (22) can be simplified to
@@
with $\mathcal{O}\left(d_{0}^{3}t^{2}e^{\lambda_{1}t}\right)\rightarrow 0$ again because $d_{0}\ll 1$. It follows that, the first term on the right hand side of Eq. (25) represents the asymptotic solution [Eq. (10)] of the variational equations [Eq. (9)] in the limit of large time, as also pointed out for example in Eq. (30) of Ref. [14], i.e. $\bm{w}_{i}(t)\approx\bm{w}_{0}e^{\lambda_{1}t}$. As a result, we write
@@
Now let us turn back to expressing $J_{k}$ [Eq. (20)] after one renormalization time $\tau$. In order to address that question, we note that for small enough values of $\tau$, we have $\lVert\bm{\delta}_{i}(\tau)\rVert\approx\lVert\bm{w}_{i}(\tau)\rVert\approx d_{0}$ leading to
@@
Consequently, we write based on Eq. (26) and Eq. (27)
@@
where we consider only leading orders. Therefore, it turns out that for $d_{0}\rightarrow 0$, the GALI as defined using the VM and the MPM practically coincide for fixed $\tau$.
Let us now look at the different numerical inaccuracies in computing the GALI in case of the MPM [Eq. (20)]. We first investigate the influence of machine precision when computing GALI after one renormalization time $\tau$, i.e.
@@
The magnitude of the initial separation vectors at the renormalization time [Eq. (21)] also impacts the computation of the GALI using the MPM. If we assume that the reference orbit $\bm{x}(t)$ is bounded (i.e. an orbit which does not blow up to infinity) then $\widetilde{\bm{x}}_{m}(t)=\bm{x}_{m}(t)+\mathcal{O}(\varepsilon)$ for all time, which also applies for all neighboring orbits $\bm{y}_{i}(t)$. As a result,
@@
with $\bm{y}_{i,m}$ being phase space coordinates of the $i^{th}$ nearby orbits from the reference orbit with element $x_{m}$. Consequently, the renormalization procedure [Eq. (21)] gives
@@
so that the leading order term in the numerical error of the GALI due to the magnitude of the separation vector reads
@@
3.2 Dependence of the GALI on the global numerical truncation errors
@@
In Ref. [29], it was argued that a numerical scheme of order $p$, with fixed integration time step $h$, affects the time propagation of the tangent vectors at time $T$ by a term $\varepsilon_{T}=\mathcal{O}\left(h^{p}\right)$. This term accounts for the global truncation error on each tangent vector at time $T$. Then, the numerical error of the GALI when including the global truncation errors can be obtained following the same steps as above, leading to
@@
Figure 1: Theoretical reliable regions for the computation of GALI of order $k$ with the (blue triangles) VM and (red squares) MPM with unit renormalization time, $\tau\approx 1$. The labels are as follows: $\nu=log_{10}\varepsilon$, $d_{0}=\lVert w_{0}\rVert=\lVert\delta_{0}\rVert$, $\sigma=\log_{10}\varepsilon_{T}$. Panel (a): $\varepsilon\lesssim\varepsilon_{T}\lesssim\varepsilon^{1/2}$. Panel (b): $\varepsilon^{1/2}\lesssim\varepsilon_{T}$.
In Fig. 1, we plot the dependence of the $\mbox{GALI}_{k}$ against $d_{0}$ at fixed $\tau=1$ value. The diagrams obtained define the reliable regions of $\mbox{GALI}_{k}$, by estimating the expressions of Eq. (36), the blue filled-triangles and Eq. (37), the red filled-circles. More specifically, we examine two cases namely when $\varepsilon<\varepsilon_{T}<\varepsilon^{1/2}$ shown in Fig. 1(a), and $\varepsilon_{T}>\varepsilon^{1/2}$ in Fig. 1(b). Note that in Fig. 1, the $\nu=\log_{10}\varepsilon$ and $\sigma=\log_{10}\varepsilon_{T}$. Overall, we see that reliable regions for the computation of GALI are independent of the GALI order. In addition the $\mbox{GALI}_{k}$ computed using the VM is independent of the magnitude of the initial tangent vectors $d_{0}$, Figs. 1(a-b). In fact, in this case the computed values of the GALI only depend on the global truncation error $\varepsilon_{T}$ of the numerical integration scheme, since the latter is general much larger than the machine precision $\varepsilon$.
On the other hand, the results of the reliable regions using the MPM can be separated in two cases. Whether $\varepsilon_{T}\in[\varepsilon,\varepsilon^{1/2}]$, we obtain a $V-$shape with minimum located at $d_{0}\sim\varepsilon^{1/2}$, red curve in Fig. 1(a). In addition, when $\varepsilon_{T}>\varepsilon$, the $V-$shape loses its pointy edge for a flat bottom, which means that the most accurate results using the MPM is also dependent on the $\varepsilon_{T}$ value, Fig. 1(b).
@@ 4.1 Hamiltonian models and computational setup
The first model of our interest is the well known two-dimensional (2D) HH system, whose Hamiltonian function reads [36, 44]
@@
with $x_{1}$ and $x_{2}$, the positions, and $p_{1}$ and $p_{2}$ their associated canonical momenta. This model gives a generic description of interactions between two asymmetric celestial bodies whose motion is restricted to a two dimensional plane (e.g. the motion of a planet around its sun, a star around the center of its galaxy, etc.) or molecules with asymmetric distributions of charges. As an additional remark, the value of the Hamiltonian function itself tunes the strength of the nonlinearity.
The second model we investigate is the multidimensional $\beta$-FPUT model which consists of a chain of $N$ coupled nonlinear oscillators. Its Hamiltonian function is expressed as [45, 37]
@@
where $x_{i}$ is the displacement of the $i^{\textit{th}}$ oscillator, $p_{i}$ its conjugate momentum, and the parameter $\beta$ controls the strength of the nonlinear terms. The $\beta$-FPUT model is used in preference to the $\alpha$-FPUT due to the absence of escaping orbits in the former, which greatly simplifies our work. The FPUT system is a generic central model in the wide and intensive field of nonlinear lattice dynamics. As such, it is involved in investigating heat conduction [46], thermalization [47, 48], ergodicity [49, 50], solitons [51], breathers [52] and chaos [53] and the possible connection between the aforementioned physical processes, see [38, 54, 55] and references therein. In our numerical simulations, we set $\beta=1$ and tune the nonlinearity of the system by changing the value of the energy $\mathcal{H}_{N}$. Further, fixed boundary conditions (i.e., $x_{0}=x_{N+1}=0$) are assumed at the two edges of the lattice.
The models above provide a perfect setup to investigate the behavior of the chaos indicators for different dynamical regimes encountered in nonlinear systems. In that regard they are extremely appreciated in computational physics, e.g. [25, 56, 26, 29, 57]. In case of the HH, an initial condition of a specific nature can be pinpointed once we generate the Poincaré surface of section (PSS) associated with a particular $\mathcal{H}$ [Eq. (38)] (a detailed description of the elaboration of a PSS can be found in Ref. [58]). Figure 2 depicts an example of such PSS defined by $x_{1}=0$ with $\mathcal{H}=0.125$.
Figure 2: 2D HH system The PSS of the HH model [Eq. (38)] at $x_{1}=0$. For each point $(x_{2},p_{2})$ within the PSS, the remaining coordinates $p_{1}$ is found such that $p_{1}>0$ with $\mathcal{H}=0.125$.
Turning now to the $\beta-$FPUT lattice model, the determination of the orbits with a specific nature is more complex mainly due to the dimensionality of the problem. Fortunately, there exist orbits which are rather simple to implement and whose vicinity has been extensively studied, the so-called simple periodic orbits of order 1 (SPO1) [59]. Thus in our study, we work with perturbations of these SPO1 orbits. In short, the SPO1 is a nonlinear continuation of the mode of the linearized system with wave number $\nu=(N+1)/2$, taking the form
@@
for $j=1,2,\ldots,\frac{N+1}{2}$ with the momentum of all particles set to $p_{j}(0)=0$. Given a finite number of particles the SPO1 orbits are regular for energy values $\mathcal{H}_{N}$ below the so-called first destabilization energy threshold $\mathcal{H}_{N}^{c}$, while being chaotic pass that threshold [59]. The $\mathcal{H}_{N}^{c}$ can be found using the standard linear stability analysis (see [13, 60, 52, 61] for details).
During all numerical integration, we are using the symplectic $ABA864$ order $4$ numerical scheme [62, 63, 64] which has proven to be efficient for evolving the tangent dynamics (using the standard method for variational equations [5, 6]) for low and high precision numerical computations [27, 64]. We perform double precision computations, i.e. $\varepsilon\approx 10^{-16}$ as it is the default numeric type of almost all modern computational tasks, and required for accurate chaotic simulations. Furthermore, we control the global truncation error $\varepsilon_{T}$ through the relative energy error $E_{r}$ defined at every time $t>0$ with
@@
for both the HH and $\beta-$FPUT models respectively. For SIs, this means changing the integration time step $h$, as this quantity remains fixed during the whole evolution of the system’s trajectory.
In addition the initial conditions of the set of orthogonal tangent vectors have components arising from a random uniform distribution with zero mean. Note that each component is duly rescaled to achieve the desired magnitude. For completeness sake in A we provide a general algorithm for the computation of GALI using the MPM. The program implemented using this algorithm computes the time evolution of the $\mbox{GALI}_{k}$ up to a given time $t=T_{M}$. Further, it can easily be amended the perform similar tasks as for the VM counterpart (see e.g. Table 2 in Ref. [15]).
4.2 Hénon-Heiles model
Figure 3: 2D HH system. Time evolution of (a-b) the two largest finite-time LEs $\lambda_{1}$ (blue) and $\lambda_{2}$ (green), and (c-d) the $\mbox{GALI}_{2}$ (blue), $\mbox{GALI}_{3}$ (green) and $\mbox{GALI}_{4}$ (red) for a regular orbit with coordinate $x_{1}=0$, $p_{1}>0$, $x_{2}=0.558$ and $p_{2}=0$ on the Poincaré surface of section in Fig. 2 of the HH with $\mathcal{H}=0.125$. The results are generated using the (a),(c) the VM and (b),(d) the MPM. The filled black lines in (a-b) guide the eye with slope $-1$, while the filled, dashed, dotted-dotted dashed ones in (c-d) have slopes 00 [horizontal lines], $-2$ and $-4$ respectively. All the calculations are performed with $d_{0}=10^{-8}$.

### v2

@@ 3.1 The variational method
In this section, we focus on evaluating the numerical uncertainty of the computation of the GALI using the variational approach. It is worth noting that the VM constitutes the sole approach considered in previous studies (see e.g. [17, 18]) for the computation of the GALI. Thus, later on, we will prove that the definition of the GALI in the variational framework is equivalent to the definition using the MPM. The variational equations are obtained by linearizing the equations of motion [Eq. (2)] around a reference orbit $\bm{x}(t)$ in the phase space. Considering a set of $k$ orthogonal deviation (tangent) vectors
@@
where $\bm{D}^{2}\mathcal{H}\left(\bm{x}(t)\right)$ is the Hessian matrix of the Hamiltonian evaluated at the reference orbit $\bm{x}(t)$, and the subscript $i$ indexes the tangent vector. As initial conditions for the variational equations, the common practice is to choose random coordinates for the deviation vectors, which are then rescaled in such a way that they all share the same norm, i.e., $\lVert\bm{w}_{i}(0)\rVert=\lVert\bm{w}_{i,0}\rVert=d_{0}$ ($d_{0}>0$).
@@
In what follows, we first assume that the trajectory of the system $\bm{x}(t)$ and its tangent dynamics $\bm{w}_{i}(t)$ are known in a closed form. Consequently, we define the GALI of order $k$ (denoted as $G_{k}$ for the VM) of the orbit $\bm{x}$ after a renormalization time $\tau$ [17, 18] as
@@
is the unit vector of the $i^{\textit{th}}$ deviation vector $\bm{w}_{i}(\tau)$.
We first investigate the error due to the machine precision (referred to as $\varepsilon$) during the computation of the $\mbox{GALI}_{k}$ after one renormalization time $\tau$, which we denote as $G_{k}(\tau,\varepsilon)$. Starting from unit deviation vectors, i.e., $d_{0}=1$, and assuming that the renormalization time $\tau$ is small, we have $\lVert\bm{w}_{i}(\tau)\rVert\approx 1$ such that
@@
with $\widetilde{u}_{i,m}$ denoting the machine representation of $u_{i,m}$. In general, we will denote by $\widetilde{Q}$ any approximation of the variable $Q$. Thus, the computer representation of the GALI of order $k$ is
@@
in such as way that the error due to the machine precision becomes
@@
In addition to the effect of the machine precision on the computation of the GALI, the magnitude of the initial deviation vectors, $d_{0}=\lVert\bm{w}_{i}(t_{0})\rVert$, also impacts the computed GALI values at each renormalization time. By denoting with $\widetilde{G}_{k}\left(\tau,\rho\right)$ the GALIk obtained by replacing $\bm{w}_{i}(t_{0})$ by $\bm{w}_{i}^{\rho}(t_{0})=\rho\cdot\lVert\bm{w}_{i}(t_{0})\rVert$, where $\rho=d_{0}/\lVert\bm{w}_{i}(t_{0})\rVert$ is the rescaling factor, we see that the time evolution of the rescaled deviation vectors $\bm{w}_{i}^{\rho}(t)$ is given by
@@
and the magnitude of the $k$ initial deviation vectors do not affect the calculations of GALI in case of the VM. In this context, we only have to take into account round off errors for the computer representation of $\bm{u}_{i}^{\rho}$, similar to what was done in Eq. (13). Thus, the numerical inaccuracy of the GALIk at the renormalization time due to the magnitude of the deviation vectors leads to
@@
3.2 The multi-particle method
@@
the unit vector of the $i^{\textit{th}}$ separation vector in the phase space. Our first task here is to show that Eq. (20) is equivalent to Eq. (11) i.e., $J_{k}(\tau)=G_{k}(\tau)$, so that the GALIs defined using the VM and the MPM are equal.
@@
to approximate the time evolution of $\bm{d}_{i}$, with $\mathcal{O}\left(\lVert\bm{d}_{i}\rVert^{3}\right)\rightarrow 0$, because $\lVert\bm{d}_{i}\rVert\ll 1$. At this point, it is convenient for us to approximate the evolution of each element $d_{i,m}$ of the separation vector with index $i$, by the dynamics of its fastest changing coordinate corresponding to the direction of maximal stretching of the tangent vectors, whose rate of growth is the MLE [7]. Thus Eq. (22) can be simplified to
@@
with $\mathcal{O}\left(d_{0}^{3}t^{2}e^{\lambda_{1}t}\right)\rightarrow 0$ again because $d_{0}\ll 1$. It follows that, the first term on the right hand side of Eq. (25) represents the asymptotic solution [Eq. (10)] of the variational equations [Eq. (9)] in the limit of large time, as also pointed out for example in Eq. (30) of Ref. [17], i.e., $\bm{w}_{i}(t)\approx\bm{w}_{0}e^{\lambda_{1}t}$. As a result, we write
@@
Now let us turn back to expressing $J_{k}$ [Eq. (20)] after one renormalization time $\tau$. In order to perform that task, we note that for small enough values of $\tau$, we have $\lVert\bm{\delta}_{i}(\tau)\rVert\approx\lVert\bm{w}_{i}(\tau)\rVert\approx d_{0}$ leading to
@@
Consequently, based on Eq. (26) and Eq. (27), we write
@@
where we consider only leading order terms. Therefore, it turns out that for $d_{0}\rightarrow 0$, the GALIk as defined using the VM and the MPM practically coincide for fixed $\tau$.
Let us now look at the different numerical inaccuracies in computing the GALIk in the case of the MPM [Eq. (20)]. We first investigate the influence of machine precision when computing the GALIk after one renormalization time $\tau$, i.e.,
@@
The magnitude of the initial separation vectors at the renormalization time [Eq. (21)] also impacts the computation of the GALI by the MPM. If we assume that the reference orbit $\bm{x}(t)$ is bounded (i.e., an orbit which does not escape to infinity) then $\widetilde{\bm{x}}_{m}(t)=\bm{x}_{m}(t)+\mathcal{O}(\varepsilon)$ for all time, which also applies for all neighboring orbits $\bm{y}_{i}(t)$. As a result,
@@
with $\bm{y}_{i,m}$ being the phase space coordinates of the $i^{th}$ nearby orbit from the reference orbit with coordinate $x_{m}$. Consequently, the renormalization procedure [Eq. (21)] gives
@@
so that the leading order term in the numerical error of the GALIk due to the magnitude of the separation vector becomes
@@
3.3 Dependence of the generalized alignment index values on the global numerical truncation errors
@@
In Ref. [31], it was argued that a numerical integration scheme of order $p$, with fixed integration time step $h$, affects the time propagation of the tangent vectors at time $T$ by a term $\varepsilon_{T}=\mathcal{O}\left(h^{p}\right)$. This term accounts for the global truncation error on each tangent vector at time $T$. Then, the numerical error of the GALIk when we include the global truncation errors can be obtained following the same steps as above, leading to
@@
In Fig. 1 we plot the dependence of the $\mbox{GALI}_{k}$ against $d_{0}$ at fixed $\tau=1$ value. The diagrams obtained define the reliable regions of $\mbox{GALI}_{k}$, by estimating the expressions of Eq. (36) (blue filled-triangles) and Eq. (37) (red filled-circles). More specifically, we examine two cases namely when $\varepsilon<\varepsilon_{T}<\varepsilon^{1/2}$ shown in Fig. 1(a), and $\varepsilon_{T}>\varepsilon^{1/2}$ in Fig. 1(b). Note that in Fig. 1 $\nu=\log_{10}\varepsilon$ and $\sigma=\log_{10}\varepsilon_{T}$. Overall, we see that reliable regions for the computation of the GALI are independent of the GALI order. In addition the $\mbox{GALI}_{k}$ computed using the VM is independent of the magnitude of the initial tangent vectors $d_{0}$. In fact, in this case the computed values of the GALI only depend on the global truncation error $\varepsilon_{T}$ of the numerical integration scheme, since the latter is, in general, much larger than the machine precision $\varepsilon$. On the other hand, the results of the reliable regions using the MPM can be separated in two cases. Whenever $\varepsilon_{T}\in[\varepsilon,\varepsilon^{1/2}]$, we obtain a $V-$shape with minimum located at $d_{0}\sim\varepsilon^{1/2}$ [red-colored shape in Fig. 1(a)]. Furthermore, when $\varepsilon_{T}>\varepsilon$, the $V-$shape loses its pointy edge to a flat bottom, which means that the most accurate results using the MPM are also dependent on the $\varepsilon_{T}$ value [Fig. 1(b)].
Figure 1: Theoretical reliable regions for the computation of the GALI of order $k$ by the VM (blue triangles) and the MPM (red circles) for unit renormalization time $\tau=1$. The labels are as follows: $\nu=log_{10}\varepsilon$, $d_{0}=\lVert w_{0}\rVert=\lVert\delta_{0}\rVert$, $\sigma=\log_{10}\varepsilon_{T}$. Panel (a): $\varepsilon\lesssim\varepsilon_{T}\lesssim\varepsilon^{1/2}$. Panel (b): $\varepsilon^{1/2}\lesssim\varepsilon_{T}$.
@@ 4.1 Hamiltonian models and computational setup
The first model we study is the well known, two degree of freedom (2D) HH system, whose Hamiltonian function reads [42]
@@
with $x_{1}$ and $x_{2}$ being the system’s generalized positions, and $p_{1}$ and $p_{2}$ their associated canonical momenta. This model gives a generic description of interactions between two asymmetric celestial bodies whose motion is restricted to a two dimensional plane (e.g., the motion of a planet around its sun, a star around the center of its galaxy, etc.) or molecules with asymmetric distributions of charges. As an additional remark we note that the value of the Hamiltonian function itself tunes the strength of the system’s nonlinearity.
The second model we investigate is the multidimensional $\beta$-FPUT model which consists of a chain of $N$ coupled nonlinear oscillators. Its Hamiltonian function is expressed as [43]
@@
where $x_{i}$ is the displacement of the $i^{\textit{th}}$ oscillator, $p_{i}$ its conjugate momentum, and the parameter $\beta$ controls the strength of the nonlinear terms. The $\beta$-FPUT model is used in preference to the $\alpha$-FPUT due to the absence of escaping orbits in the former, which greatly simplifies our work. The FPUT system is a generic central model in the wide and intensive field of nonlinear lattice dynamics. As such, it has been considered in investigations of heat conduction [49], thermalization [50], ergodicity [51], solitons [52], breathers [53], classical and quantum chaos [54, 55] and the possible connection between the aforementioned physical processes, see for example [44] and references therein. In our numerical simulations, we set $\beta=1$ and tune the nonlinearity of the system by changing the value $\mathcal{H}_{N}$ of the energy. Furthermore, fixed boundary conditions (i.e., $x_{0}=x_{N+1}=0$) are imposed at the two edges of the lattice.
These two models provide a perfect setup to investigate the behavior of chaos indicators for different dynamical regimes encountered in nonlinear systems. In that regard they are extremely appreciated in computational physics, see e.g. [27, 28, 31]. In the case of the HH, an initial condition of a specific nature can be identified once we generate the system’s Poincaré surface of section (PSS) associated with a particular $\mathcal{H}$ [Eq. (38)]. Figure 2 depicts an example of a PSS defined by $x_{1}=0$ and $p_{1}>0$ with $\mathcal{H}=0.125$.
Figure 2: 2D HH system. The PSS of the HH model [Eq. (38)] at $x_{1}=0$ for $\mathcal{H}=0.125$. For each point $(x_{2},p_{2})$ of the PSS, the remaining coordinate $p_{1}$ is found such that $p_{1}>0$.
Turning now to the $\beta$-FPUT lattice model, the determination of the orbits with a specific dynamical nature is more complex mainly due to the high dimensionality of the system’s phase space. Fortunately, there exist orbits which are rather simple to locate and whose vicinity has been extensively studied, the so-called simple periodic orbits of order 1 (SPO1) [56, 57]. Thus in our study, we work with perturbations of these SPO1 orbits. In short, the SPO1 is a nonlinear continuation of the mode of the linearized system with wave number $\nu=(N+1)/2$, taking the form
@@
for $j=1,2,\ldots,\frac{N+1}{2}$ with the momentum of all particles set to $p_{j}(0)=0$. Given a finite number of oscillators the SPO1 orbits are regular for energy values $\mathcal{H}_{N}$ below the so-called first destabilization energy threshold $\mathcal{H}_{N}^{c}$, while being chaotic past that threshold [57]. The $\mathcal{H}_{N}^{c}$ can be found using the standard linear stability analysis (see [16, 58, 53, 59] for details).
Let us now explain how we practically compute the LEs and the GALIs. To estimate the first $k$ largest LEs, we integrate simultaneously the equations of motion along with a set of $k$ variational equations or nearby-orbits. For each simulation, we start from $k$ random and linearly independent tangent vectors [$\bm{d}_{i}(0)$ or $\bm{w}_{i}(0)$] which are rescaled to achieve the size $d_{0}$. Then, the LEs are estimated using the ‘standard method’ [5, 6, 7], leaving the renormalization time $\tau$, and the deviation vector size $d_{0}$, as free parameters. A pseudo-code for the implementation of VM for this approach can be found in [7]. For the sake of completeness, in A we also provide a pseudo-code for the computation of the GALI using the MPM, which is based on the approach developed in [60] (see Ref. [18] for the VM counterpart).
We evolve the system’s phase space orbits and the associated neighboring trajectories using the symplectic integration scheme $ABA864$ of order $4$ [61, 62, 29, 63] with fixed time step, $h$. The value of $h$ controls the global truncation error $\varepsilon_{T}$ which we track at every time, $t>0$, using the relative energy error
@@
for the HH and the $\beta$-FPUT models respectively. In the case of the VM the tested orbit and the related variational equations are simultaneously integrated by the $ABA864$ integrator using the so-called ‘tangent map method’ [27, 28]. In addition, we perform double precision computations, i.e., $\varepsilon\approx 10^{-16}$, as this is the default numeric type of almost all modern computational tasks.
4.2 The Hénon-Heiles model